What is the present value of a single cash flow of $25,000 received at the end of 10 years, if we assume a discount rate of 5% annually? With a discount rate of 7%
1
Expert's answer
2018-04-18T06:05:07-0400
If we assume a discount rate of 5% annually, then the present value of a single cash flow of $25,000 received at the end of 10 years is:
PV = CF/(1 + r)^t = 25,000/1.05^10 = $15,347.83.
With a discount rate of 7%:
PV = CF/(1 + r)^t = 25,000/1.07^10 = $12,708.73.